OpenStudy (anonymous):

find the range of x*x -5x+6>0

hartnn (hartnn):

do you know what is the product roots formula for \(ax^2+bx+c =0\) ?

hartnn (hartnn):

**product of roots

OpenStudy (anonymous):

c/a

hartnn (hartnn):

for \(ax^2+bx+c =0\) product of roots = \(\large \dfrac{c}{a}\) so for \(5x^2+14x+k=0\) what will be the product of roots ?

hartnn (hartnn):

identify c and a for 5x^2 +14x+k

OpenStudy (anonymous):

k/5

hartnn (hartnn):

correct! now since , one root is reciprocal of other, their product will be 1 !!

hartnn (hartnn):

so, get k from k/5 =1 piece of cake :P

OpenStudy (anonymous):

what is meaning of reciprocal rootd dude

hartnn (hartnn):

ohh! reciprocal means if one root is a, then other is 1/a ! their product = a*(1/a) =1

OpenStudy (anonymous):

so,here they didn't tell that product of the roots are reciprocal

hartnn (hartnn):

no, roots are reciprocal like a and 1/a from that we got that product of roots must be 1 hence we took, c/a = k/5 = 1

hartnn (hartnn):

makes sense ?

OpenStudy (anonymous):

ya.i got it.thank u.

OpenStudy (anonymous):

may iknow from which country u r?

hartnn (hartnn):

welcome ^_^ I am Indian :)

hartnn (hartnn):

Since you're new here, \(\Huge \mathcal{\text{Welcome To OpenStudy}\ddot\smile} \)

OpenStudy (anonymous):

state?

OpenStudy (anonymous):

this is the question from oracle interview

hartnn (hartnn):

oh! best of luck for your preparations for the interview :)

OpenStudy (anonymous):

if first term of AP is 5 .16th term is 45.find the sum of first 16 terms.

hartnn (hartnn):

know the formula for n'th term of AP and sum of terms in AP ?

OpenStudy (anonymous):

a+(n-1)d

OpenStudy (anonymous):

n(n+1)/2

hartnn (hartnn):

\(\Large S_{16} = \dfrac{16}{2}(a_1+a_{16})\)

hartnn (hartnn):

general formula is \(\large S_n = (n/2)(a_1 +a_n)\)

OpenStudy (anonymous):

how u got this one

hartnn (hartnn):

There are 2 formulas for the sum, the one i posted and other is : \(\large S_n = (n/2)(2a+(n-1)d)\)

hartnn (hartnn):

since we know first and last term, using first formula is much easier and faster! :)

OpenStudy (anonymous):

this one i know

OpenStudy (anonymous):

i studied second one in r.s.agarwal

hartnn (hartnn):

both formulas are equally important :)

OpenStudy (anonymous):

from which state u r?

hartnn (hartnn):

i messaged you

hartnn (hartnn):

It not good to disclose personal information on the internet ...

hartnn (hartnn):

*Its

OpenStudy (anonymous):

just state only i asked u.if u have any problem leave it

hartnn (hartnn):

ok :)

OpenStudy (anonymous):

is it possible to find using second formula

hartnn (hartnn):

ofcourse! but it would be a longer method.

OpenStudy (anonymous):

there 15d=40

hartnn (hartnn):

correct!

OpenStudy (anonymous):

then what is the next step

hartnn (hartnn):

you have a1, and d, and n whats stopping you from using the formula you know!

OpenStudy (anonymous):

ok

hartnn (hartnn):

15 d =40 d= 40/15 = 8/3

OpenStudy (anonymous):

find the range of x*x -5x+6>0

hartnn (hartnn):

ask new question in new post :)

OpenStudy (anonymous):

find the range of x*x -5x+6>0